OpenStudy (anonymous):

In ABC /C is a right angle. Find the remaining sides and angles. Round your answers to the nearest tenth. A=3, C =19  

OpenStudy (anonymous):

a. M/A = 9.1°,m/B = 80.9°, b = 18.8 c.M/A = 14.5°,m/B = 75.5°, b = 18.8 b.M/A = 80.9°, m/B = 9.1°, b = 18.8 d.M/A = 75.5°,m/B = 14.5°, b = 18.8